# On the Convergence Analysis of Muon

### v1

@@ Appendix B Nonconvex
Moreover, when $t>0$, we can note that
@@
Denote $\Delta=f(W_{0})-f^{*}$ and set $\eta=\sqrt{\frac{(1-\beta)\Delta}{rTL}}$, we can get
@@
Denote $\Delta=f(W_{0})-f^{*}$ and set $\eta=\sqrt{\frac{(1-\beta)\Delta}{TL_{*}}}$, we can get
@@
Denote $d=mn$, $J_{t}=\mathsf{vec}(U_{t}V_{t}^{\top})^{\top}H_{t}\mathsf{vec}(U_{t}V_{t}^{\top})$, $H_{t}=\nabla^{2}f_{v}(\mathsf{vec}(W_{t}))$, and $f_{v}(\mathsf{vec}(W_{t}))=f(W_{t})$. Set $\eta_{t}=\eta$. Taking the Taylor expansion at $W_{t}$, we have
$f(W_{t+1})-f(W_{t})=$ $\langle\nabla f(W_{t}),W_{t+1}-W_{t}\rangle+\frac{1}{2}\mathsf{vec}(W_{t+1}-W_{t})^{\top}\nabla^{2}f_{v}(\mathsf{vec}(W_{t}))\mathsf{vec}(W_{t+1}-W_{t})$
$+\frac{1}{6}\sum_{i,j,k=1}^{d}[\nabla^{3}f_{v}(\theta)]_{ijk}\mathsf{vec}(W_{t+1}-W_{t})_{i}\mathsf{vec}(W_{t+1}-W_{t})_{j}\mathsf{vec}(W_{t+1}-W_{t})_{k}$
$\leq$ $-\eta\langle\nabla f(W_{t}),U_{t}V_{t}^{\top}\rangle+\frac{\eta^{2}J_{t}}{2}+\frac{s\eta^{3}\|\mathsf{vec}(U_{t}V_{t}^{\top})\|_{\rm F}^{3}}{6}$
@@
where $\theta\in\mathbb{R}^{d}$ can be some vectors between $\mathsf{vec}(W_{t})$ and $\mathsf{vec}(W_{t+1})$, and the first inequality is due to 4.7.
Denote $J=\frac{1}{T}\sum_{t=0}^{T-1}J_{t}$. Summing over $t=0,1,...,T-1$, we can get
@@
Denote $\Delta=f(W_{0})-f^{*}$. In experiments, we found that usually $J>0$. When $J>0$ and set $\eta=\sqrt{\frac{2\Delta}{JT}}$, we can get
@@
Thus, when $J>0$, we can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(J\Delta\epsilon^{-2})$.
@@
Thus, when $J\leq 0$, we can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(s^{1/2}r^{3/4}\Delta\epsilon^{-3/2})$. ∎
@@ Appendix C Star-Convex
In this section, we first present the proof of Theorem 4.11. Then, we present an additional convergence analysis of Muon under 3.1 and 4.7 in Section C.3 and Section C.2 respectively.
@@
where we apply the Cauchy–Schwarz inequality in the second inequality, and in the third inequality we use 4.10.
Set $\eta=\eta$ and combine (13) and (14). We have
@@
Thus, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, we can take $\eta=O(\frac{\epsilon}{L_{*}D_{{\rm op}}})$, and the complexity is $O(L_{*}D_{{\rm op}}^{2}\epsilon^{-1}\text{log}\frac{\Delta}{\epsilon})$, where $\Delta=f(W_{0})-f^{*}$.
@@
C.2 Convergence analysis with 4.7
Theorem C.1.
Under 4.9, 4.7 and 4.10, if we apply Algorithm 2 with $\eta_{t}=\eta$, we have
$f(W_{T})-f^{*}\leq$ $\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T}(f(W_{0})-f^{*})+\sum_{t=0}^{T-1}\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T-1-t}\frac{\eta^{2}J_{t}}{2}+\frac{s\eta^{2}D_{\rm op}r^{3/2}}{6}.$
where $J_{t}=\mathsf{vec}(U_{t}V_{t}^{\top})^{\top}H_{t}\mathsf{vec}(U_{t}V_{t}^{\top})$, $H_{t}=\nabla^{2}f_{v}(\mathsf{vec}(W_{t}))$, and $f_{v}(\mathsf{vec}(W_{t}))=f(W_{t})$.
Denote $\Delta=f(W_{0})-f^{*}$ and $\hat{J}=\frac{1}{T}\sum_{t=0}^{T-1}|J_{t}|$. We have
$f(W_{T})-f^{*}$ $\leq\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T}\Delta+\frac{\eta^{2}\hat{J}T}{2}+\frac{s\eta^{2}D_{\rm op}r^{3/2}}{6}.$
When $\hat{J}>0$, we can set $\eta=\min\left\{\frac{D_{{\rm op}}}{T}\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\hat{J}}\right),D_{\rm op}\right\}$. We have
$f(W_{T})-f^{*}\leq$ $\frac{D_{{\rm op}}^{2}\hat{J}}{T}+\frac{D_{{\rm op}}^{2}\hat{J}}{2T}\left[\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\hat{J}}\right)\right]^{2}+\frac{sr^{3/2}D_{{\rm op}}^{3}}{6T^{2}}\left[\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\hat{J}}\right)\right]^{2}$
$\leq$ $\tilde{O}\left(\frac{D_{{\rm op}}^{2}\hat{J}}{T}\right).$
When $\hat{J}=0$, we can set $\eta=\min\left\{\frac{D_{{\rm op}}}{T}\log\left(\frac{T^{2}\Delta}{D_{{\rm op}}^{3}sr^{3/2}}\right),D_{\rm op}\right\}$. We have
$\leq$ $\tilde{O}\left(\frac{sr^{3/2}D_{{\rm op}}^{3}}{T^{2}}\right).$
@@
Denote $\Delta=f(W_{0})-f^{*}$ and $\tilde{J}=\sum_{t=0}^{T-1}\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T-1-t}\frac{\eta}{D_{\rm op}}J_{t}$. We have
$f(W_{T})-f^{*}$ $\leq\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T}\Delta+\frac{\eta D_{\rm op}\tilde{J}}{2}+\frac{s\eta^{2}D_{\rm op}r^{3/2}}{6}.$
Thus, the convergence complexity can be depending on $\tilde{J}$, which can be viewed as a weighted average of $J_{t}$.
In experiments, we found that usually $J_{t}>0$. Note that if $f$ is strict convex and if $U_{t}V_{t}^{\top}\neq 0$, then $J_{t}$ is strictly larger than 0. Thus, to get a more illustrative and clearer result, if we denote $\hat{J}=\frac{1}{T}\sum_{t=0}^{T-1}|J_{t}|$, then we have
$f(W_{T})-f^{*}$ $\leq\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T}\Delta+\frac{\eta^{2}\hat{J}T}{2}+\frac{s\eta^{2}D_{\rm op}r^{3/2}}{6}.$
When $\hat{J}>0$, we can set $\eta=\min\left\{\frac{D_{{\rm op}}}{T}\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\hat{J}}\right),D_{\rm op}\right\}$. We have
$f(W_{T})-f^{*}\leq$ $\frac{D_{{\rm op}}^{2}\hat{J}}{T}+\frac{D_{{\rm op}}^{2}\hat{J}}{2T}\left[\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\hat{J}}\right)\right]^{2}+\frac{sr^{3/2}D_{{\rm op}}^{3}}{6T^{2}}\left[\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\hat{J}}\right)\right]^{2}$
$\leq$ $\tilde{O}\left(\frac{D_{{\rm op}}^{2}\hat{J}}{T}\right).$
When $\hat{J}=0$, we can set $\eta=\min\left\{\frac{D_{{\rm op}}}{T}\log\left(\frac{T^{2}\Delta}{D_{{\rm op}}^{3}sr^{3/2}}\right),D_{\rm op}\right\}$. We have

### v2

@@ Appendix B Nonconvex
Moreover, when $t>0$, we note that
@@
Set $\eta=\sqrt{\frac{(1-\beta)\Delta}{rTL}}$, we can get
@@
Set $\eta=\sqrt{\frac{(1-\beta)\Delta}{TL_{*}}}$, we can get
@@
Denote $J_{t}=\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})^{\top}H_{t}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})$, $H_{t}=\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))$, and $f_{v}(\overline{\mathrm{vec}}(W_{t}))=f(W_{t})$. Set $\eta_{t}=\eta$. Taking the Taylor expansion at $W_{t}$, we have
$f(W_{t+1})-f(W_{t})=$ $\langle\nabla f(W_{t}),W_{t+1}-W_{t}\rangle+\frac{1}{2}\overline{\mathrm{vec}}(W_{t+1}-W_{t})^{\top}\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))\overline{\mathrm{vec}}(W_{t+1}-W_{t})$
$+\frac{1}{6}\sum_{i,j,k=1}^{d}[\nabla^{3}f_{v}(\theta)]_{ijk}\overline{\mathrm{vec}}(W_{t+1}-W_{t})_{i}\overline{\mathrm{vec}}(W_{t+1}-W_{t})_{j}\overline{\mathrm{vec}}(W_{t+1}-W_{t})_{k}$
$\leq$ $-\eta\langle\nabla f(W_{t}),U_{t}V_{t}^{\top}\rangle+\frac{\eta^{2}J_{t}}{2}+\frac{s\eta^{3}\|\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})\|_{\rm F}^{3}}{6}$
@@
where $\theta\in\mathbb{R}^{mn}$ can be some vectors between $\overline{\mathrm{vec}}(W_{t})$ and $\overline{\mathrm{vec}}(W_{t+1})$, and the first inequality is due to 4.13.
Denote $J=\frac{1}{T}\sum_{t=0}^{T-1}J_{t}$. Summing over $t=0,1,\dots,T-1$, we can get
@@
In experiments, we found that usually $J>0$. When $J>0$ and $\eta=\sqrt{\frac{2\Delta}{JT}}$, we can get
@@
Thus, it is possible that Muon can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(J\Delta\epsilon^{-2})$.
@@
Thus, when $J\leq 0$, Muon can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(s^{1/2}r^{3/4}\Delta\epsilon^{-3/2})$.
B.4 Convergence analysis of Algorithm 1 with 4.13 in the stochastic setting
Theorem B.1.
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(W_{t})\|_{*}]\leq\frac{\Delta}{T\eta}+\frac{J^{*}\eta}{2}+\frac{2\sigma\sqrt{r(1-\beta)}}{\sqrt{(1+\beta)B}}+\frac{2\beta\sigma\sqrt{r}}{(1-\beta)T\sqrt{B}}+\frac{2\eta\beta J^{*}}{1-\beta}+O\left(\frac{sr^{3/2}\eta^{2}}{1-\beta}\right).$
where $J^{*}=\max\{\mathbb{E}[\hat{J}],\mathbb{E}[J]\}$, $\hat{J}=\frac{1}{T}\sum_{t=0}^{T-2}\hat{J}_{t}$, $J=\frac{1}{T}\sum_{t=0}^{T-1}J_{t}$, $J_{t}=\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})^{\top}\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})$, $\hat{J}_{t}=\overline{\mathrm{vec}}(U_{g,t}V_{g,t}^{\top})^{\top}\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})$, $H_{t}=\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))$, $f_{v}(\overline{\mathrm{vec}}(W_{t}))=f(W_{t})$, and $U_{g,t},V_{g,t}$ are the the matrices of left and right singular vectors of $\nabla f(W_{t})-\nabla f(W_{t+1})$, i.e. $\nabla f(W_{t})-\nabla f(W_{t+1})=U_{g,t}S_{g,t}V_{g,t}^{\top}$.
Proof.
Denote $d=mn$, $J_{t}=\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})^{\top}H_{t}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})$, $H_{t}=\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))$, and $f_{v}(\overline{\mathrm{vec}}(W_{t}))=f(W_{t})$. Set $\eta_{t}=\eta$. Taking the Taylor expansion at $W_{t}$, we have
$\mathbb{E}[f(W_{t+1})-f(W_{t})]$
$=$ $\mathbb{E}[\langle\nabla f(W_{t}),W_{t+1}-W_{t}\rangle+\frac{1}{2}\overline{\mathrm{vec}}(W_{t+1}-W_{t})^{\top}\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))\overline{\mathrm{vec}}(W_{t+1}-W_{t})$
$+\frac{1}{6}\sum_{i,j,k=1}^{d}[\nabla^{3}f_{v}(\theta)]_{ijk}\overline{\mathrm{vec}}(W_{t+1}-W_{t})_{i}\overline{\mathrm{vec}}(W_{t+1}-W_{t})_{j}\overline{\mathrm{vec}}(W_{t+1}-W_{t})_{k}]$
$\leq$ $\mathbb{E}\left[-\eta\langle\nabla f(W_{t}),U_{t}V_{t}^{\top}\rangle+\frac{\eta^{2}J_{t}}{2}+\frac{s\eta^{3}\|\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})\|_{F}^{3}}{6}\right]$
$\leq$ $\mathbb{E}\left[-\eta\langle M_{t},U_{t}V_{t}^{\top}\rangle+\eta\langle M_{t}-\nabla f(W_{t}),U_{t}V_{t}^{\top}\rangle+\frac{\eta^{2}J_{t}}{2}\right]+\frac{s\eta^{3}r^{3/2}}{6}$
$\leq$ $\mathbb{E}\left[-\eta\|M_{t}\|_{*}+\eta\|M_{t}-\nabla f(W_{t})\|_{*}\|U_{t}V_{t}^{\top}\|_{\text{op}}+\frac{\eta^{2}J_{t}}{2}\right]+\frac{s\eta^{3}r^{3/2}}{6}$
$\leq$ $\mathbb{E}\left[-\eta\|\nabla f(W_{t})\|_{*}+2\eta\|M_{t}-\nabla f(W_{t})\|_{*}+\frac{\eta^{2}J_{t}}{2}\right]+\frac{s\eta^{3}r^{3/2}}{6}$
where $\theta\in\mathbb{R}^{d}$ can be some vectors between $\overline{\mathrm{vec}}(W_{t})$ and $\overline{\mathrm{vec}}(W_{t+1})$, the first inequality is due to 4.13.
$\mathbb{E}[\|C_{t}-M_{t}\|_{*}]\leq\sqrt{r}\mathbb{E}[\|C_{t}-M_{t}\|_{F}]\leq$ $\sqrt{\frac{1-\beta}{1+\beta}}\frac{\sigma\sqrt{r}}{\sqrt{B}}+\beta^{t}\frac{\sigma\sqrt{r}}{\sqrt{B}}$
Suppose the SVD of $\nabla f(W_{t})-\nabla f(W_{t+1})$ is $\nabla f(W_{t})-\nabla f(W_{t+1})=U_{g,t}S_{g,t}V_{g,t}^{\top}$. Then, we have
$\|\nabla f(W_{t})-\nabla f(W_{t+1})\|_{*}$
$=$ $\|U_{g,t}^{\top}[\nabla f(W_{t})-\nabla f(W_{t+1})]V_{g,t}\|_{*}$
$=$ $\text{tr}\left(U_{g,t}^{\top}[\nabla f(W_{t})-\nabla f(W_{t+1})]V_{g,t}\right)$
$=$ $\text{tr}\left((U_{g,t}V_{g,t}^{\top})^{\top}[\nabla f(W_{t})-\nabla f(W_{t+1})]\right)$
For any $i\in[d]$, taking the Taylor expansion of $\overline{\mathrm{vec}}(\nabla f(W_{t+1}))_{i}$ at $W_{t}$, we have
$[\overline{\mathrm{vec}}(\nabla f(W_{t})-\nabla f(W_{t+1}))]_{i}$
$=$ $\sum_{j=1}^{d}[\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))]_{ij}\overline{\mathrm{vec}}(W_{t}-W_{t+})_{j}+\frac{1}{2}\sum_{j,k=1}^{d}[\nabla^{3}f_{v}(\theta_{i})]_{ijk}\overline{\mathrm{vec}}(W_{t}-W_{t+1})_{j}\overline{\mathrm{vec}}(W_{t}-W_{t+1})_{k}$
$=$ $\eta\sum_{j=1}^{d}[\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))]_{ij}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})_{j}+\frac{\eta^{2}}{2}\sum_{j,k=1}^{d}[\nabla^{3}f_{v}(\theta_{i})]_{ijk}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})_{j}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})_{k}$
where $\theta_{i}\in\mathbb{R}^{d}$ can be some vectors between $\overline{\mathrm{vec}}(W_{t})$ and $\overline{\mathrm{vec}}(W_{t+1})$.
Denote $\hat{J}_{t}=\overline{\mathrm{vec}}(U_{g,t}V_{g,t}^{\top})^{\top}\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})$ , we have
$\|\nabla f(W_{t+1})-\nabla f(W_{t})\|_{*}$
$=$ $\text{tr}\left((U_{g,t}V_{g,t}^{\top})^{\top}[\nabla f(W_{t})-\nabla f(W_{t+1})]\right)$
$=$ $\eta\sum_{i,j=1}^{d}\overline{\mathrm{vec}}(U_{g,t}V_{g,t}^{\top})_{i}\nabla^{2}f_{v}(\overline{\mathrm{vec}}(W_{t}))]_{ij}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})_{j}$
$+\frac{\eta^{2}}{2}\sum_{i,j,k=1}^{d}[\nabla^{3}f_{v}(\theta_{i})]_{ijk}\overline{\mathrm{vec}}(U_{g,t}V_{g,t}^{\top})_{i}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})_{j}\overline{\mathrm{vec}}(U_{t}V_{t}^{\top})_{k}$
$\leq$ $\eta\hat{J}_{t}+\frac{s\eta^{2}r^{3/2}}{2}$
where the last inequality is due to 4.13.
Then, for $t>0$, we have
$\mathbb{E}[\|\nabla f(W_{t})-C_{t}\|_{*}]$
$=$ $\mathbb{E}[\beta\|\nabla f(W_{t})-C_{t-1}\|_{*}]$
$\leq$ $\mathbb{E}[\beta\|\nabla f(W_{t-1})-C_{t-1}\|_{*}+\beta\eta\hat{J}_{t-1}+\frac{s\eta^{2}r^{3/2}\beta}{2}]$
$\leq$ $\sum_{i=0}^{t-1}\beta^{t-i}\eta\mathbb{E}[\hat{J}_{i}]+\frac{s\eta^{2}r^{3/2}\beta}{2(1-\beta)}$
Thus, we have
$\mathbb{E}[\|\nabla f(W_{t})-M_{t}\|_{*}]\leq$ $\mathbb{E}[\|C_{t}-M_{t}\|_{*}+\|\nabla f(W_{t})-C_{t}\|_{*}]$
$\leq$ $\sqrt{\frac{1-\beta}{1+\beta}}\frac{\sigma\sqrt{r}}{\sqrt{B}}+\beta^{t}\frac{\sigma\sqrt{r}}{\sqrt{B}}+\sum_{i=0}^{t-1}\beta^{t-i}\eta\mathbb{E}[\hat{J}_{i}]+\frac{s\eta^{2}r^{3/2}\beta}{2(1-\beta)},$
and
$\mathbb{E}[f(W_{t})-f(W_{t+1})]$
$\geq$ $\mathbb{E}[\eta\|\nabla f(W_{t})\|_{*}-\frac{J_{t}}{2}\eta^{2}-2\eta\|\nabla f(W_{t})-M_{t}\|_{*}]-\frac{s\eta^{3}r^{3/2}}{6}$
$\geq$ $\mathbb{E}[\eta\|\nabla f(W_{t})\|_{*}-\frac{J_{t}}{2}\eta^{2}-\sum_{i=0}^{t-1}2\beta^{t-i}\eta^{2}\hat{J}_{i}]-\frac{s\eta^{3}r^{3/2}}{6}-2\eta\sqrt{\frac{1-\beta}{1+\beta}}\frac{\sigma\sqrt{r}}{\sqrt{B}}-2\eta\beta^{t}\frac{\sigma\sqrt{r}}{\sqrt{B}}-\frac{s\eta^{3}r^{3/2}\beta}{1-\beta}.$
Denote $\hat{J}=\frac{1}{T}\sum_{t=0}^{T-2}\hat{J}_{t}$, $J=\frac{1}{T}\sum_{t=0}^{T-1}J_{t}$ Summing over $t=0,1,\dots,T-1$, we can get
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(W_{t})\|_{*}]\leq$ $\frac{1}{T\eta}\mathbb{E}[f(W_{0})-f(W_{T})]+\frac{\eta\sum_{t=0}^{T-1}\mathbb{E}[J_{t}]}{2T}+\frac{2\sigma\sqrt{r(1-\beta)}}{\sqrt{(1+\beta)B}}+\frac{2\beta\sigma\sqrt{r}}{(1-\beta)T\sqrt{B}}$
$+2\frac{1}{T}\sum_{t=0}^{T-1}\sum_{i=0}^{t-1}2\beta^{t-i}\eta\mathbb{E}[\hat{J}_{i}]+O(sr^{3/2}\eta^{2}/(1-\beta))$
$\leq$ $\frac{1}{T\eta}\mathbb{E}[f(W_{0})-f(W_{T})]+\frac{\eta\mathbb{E}[J]}{2}+\frac{2\sigma\sqrt{r(1-\beta)}}{\sqrt{(1+\beta)B}}+\frac{2\beta\sigma\sqrt{r}}{(1-\beta)T\sqrt{B}}$
$+\frac{2\eta\beta\mathbb{E}[\hat{J}]}{1-\beta}+O(sr^{3/2}\eta^{2}/(1-\beta))$
Denote $J^{*}=\max\{\mathbb{E}[\hat{J}],\mathbb{E}[J]\}$, we can get
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(W_{t})\|_{*}]\leq\frac{\Delta}{T\eta}+\frac{J^{*}\eta}{2}+\frac{2\sigma\sqrt{r(1-\beta)}}{\sqrt{(1+\beta)B}}+\frac{2\beta\sigma\sqrt{r}}{(1-\beta)T\sqrt{B}}+\frac{2\eta\beta J^{*}}{1-\beta}+O\left(\frac{sr^{3/2}\eta^{2}}{1-\beta}\right).$
When $B=1$, $J^{*}>0$ and $\eta=\sqrt{\frac{(1-\beta)\Delta}{TJ^{*}}}$, $1-\beta=\min\{\frac{\sqrt{J^{*}\Delta}}{\sigma\sqrt{rT}},1\}$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(W_{t})\|_{*}]\leq O\left(\sqrt[4]{\frac{rJ^{*}\Delta\sigma^{2}}{T}}+\sqrt{\frac{J^{*}\Delta}{T}}+\frac{r\sigma^{2}}{\sqrt{J^{*}\Delta T}}+\frac{sr^{3/2}\Delta}{TJ^{*}}\right).$
Thus, it is possible that Muon can find an $\epsilon$-nuclear norm stationary point of $f$ with a complexity of $O(rJ^{*}\sigma^{2}\Delta\epsilon^{-4})$.
Note that $\hat{J}$ satisfies properties analogous to those of $J$ discussed in Section 4.3.
For example, if we make the same assumptions as those in Section 4.3, i.e. $H_{t}$ can be expressed as a Kronecker product $H_{t}=P_{t}\otimes Q_{t}$, where $P_{t}\in\mathbb{R}^{m\times m}$ and $Q_{t}\in\mathbb{R}^{n\times n}$. Let $\sigma_{p,t,1}\geq\sigma_{p,t,2}\geq\dots\geq\sigma_{p,t,m}$ denote the singular values of $P_{t}$, and $\sigma_{q,t,1}\geq\sigma_{q,t,2}\geq\dots\geq\sigma_{q,t,n}$ denote the singular values of $Q_{t}$. Then, we have $\hat{J}_{t}=\langle U_{g,t}^{\top}P_{t}U_{t},V_{g,t}^{\top}Q_{t}V_{t}\rangle\leq\sum_{i=1}^{r_{t}}\sigma_{p,t,i}\sigma_{q,t,i}$ (Von Neumann’s trace inequality).
Note that in the stochastic setting, the complexity of SGD for $\epsilon$-nuclear norm stationary point is $O(r^{2}L\sigma^{2}\Delta\epsilon^{-4})$. Thus, Similar to the discussions in Section 4.3, when $H_{t}$ can be represented by $P_{t}\otimes Q_{t}$ and $Q_{t}$, $P_{t}$ are relatively low-rank such that $\sum_{i=1}^{r_{t}}\sigma_{p,t,i}\sigma_{q,t,i}\ll r\sigma_{p,t,1}\sigma_{q,t,1}$, then $J^{*}\leq\max\{\hat{J},J\}\leq\max_{t}\sum_{i=1}^{r_{t}}\sigma_{p,t,i}\sigma_{q,t,i}\ll r\max_{t}L_{t}\leq rL$, and the convergence rate of Muon can be better than GD’s. (Or one of $Q_{t}$, $P_{t}$ is relatively low-rank, i.e. if $Q_{t}$ is relatively low-rank such that $\|Q_{t}\|_{{\rm op}}\approx\|Q_{t}\|\ll r\|Q_{t}\|_{{\rm op}}$, we can also have $J^{*}\leq\max_{t}\|P_{t}\|_{{\rm op}}\|Q_{t}\|_{*}\ll r\max_{t}L_{t}\leq rL$.)
@@ Appendix C Star-Convex
Proof.
Constant stepsize
First, by the $L$ Frobenius norm Lipschitz smoothness of $f$, we have
$\leq f(W_{t})-\eta_{t}\|\nabla f(W_{t})\|_{*}+\frac{rL\eta_{t}^{2}}{2}$ (16)
Then, according to (20), we have
$-\|\nabla f(W_{t})\|_{*}$ $\leq-\frac{f(W_{t})-f^{*}}{D_{{\rm op}}}$
Set $\eta_{t}=\eta$ and combine (16) and (20). We have
and
Thus, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, we can take $\eta=O(\frac{\epsilon}{rLD_{{\rm op}}})$, and the complexity is $O(rLD_{{\rm op}}^{2}\epsilon^{-1}\text{log}\frac{\Delta}{\epsilon})$, where $\Delta=f(W_{0})-f^{*}$. ∎
C.2 Proof of Theorem 4.6
Proof.
Adaptive stepsize
$\leq f(W_{t})-\eta_{t}\|\nabla f(W_{t})\|_{*}+\frac{rL\eta_{t}^{2}}{2}$
$=f(W_{t})-\frac{\|\nabla f(W_{t})\|_{*}^{2}}{2rL}.$ (17)
From the star-convex property, we have
Combining Equation 17 and Equation 18, we obtain
Then we have
where we use $\frac{\Delta_{t}}{\Delta_{t-1}}\geq 1$.
After rearranging terms, we obtain
C.3 Proof of Theorem 4.9
@@
where we apply the Cauchy–Schwarz inequality in the second inequality, and in the third inequality we use 4.4.
Set $\eta_{t}=\eta$ and combine (19) and (20). We have
@@
Thus, to reach the precision $f(W_{T})-f^{*}\leq\epsilon$, we can take $\eta=O(\frac{\epsilon}{L_{*}D_{{\rm op}}})$, and the complexity is $O(L_{*}D_{{\rm op}}^{2}\epsilon^{-1}\text{log}\frac{\Delta}{\epsilon})$, where $\Delta=f(W_{0})-f^{*}$. ∎
C.4 Proof of Theorem 4.10
Proof.
@@
C.5 Proof of Theorem 4.15
@@
Denote $\tilde{J}=\frac{1}{T}\sum_{t=0}^{T-1}\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T-1-t}J_{t}$. We have
$f(W_{T})-f^{*}$ $\leq\left(1-\frac{\eta}{D_{{\rm op}}}\right)^{T}\Delta+\frac{\eta^{2}\tilde{J}T}{2}+\frac{s\eta^{2}D_{\rm op}r^{3/2}}{6}.$
Thus, the convergence complexity of Muon can be depending on $\tilde{J}$, which can be viewed as a weighted average of $J_{t}$.
In experiments, we found that usually $J_{t}>0$. Note that if $f$ is strict convex and if $U_{t}V_{t}^{\top}\neq 0$, then $J_{t}$ is strictly larger than 0.
If $\tilde{J}>0$ and $\eta=\min\left\{\frac{D_{{\rm op}}}{T}\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\tilde{J}}\right),D_{\rm op}\right\}$. We have
$f(W_{T})-f^{*}\leq$ $\frac{D_{{\rm op}}^{2}\tilde{J}}{T}+\frac{D_{{\rm op}}^{2}\tilde{J}}{2T}\left[\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\tilde{J}}\right)\right]^{2}+\frac{sr^{3/2}D_{{\rm op}}^{3}}{6T^{2}}\left[\log\left(\frac{T\Delta}{D_{{\rm op}}^{2}\tilde{J}}\right)\right]^{2}$
$\leq$ $\tilde{O}\left(\frac{D_{{\rm op}}^{2}\tilde{J}}{T}\right).$
When $\tilde{J}\leq 0$, we can set $\eta=\min\left\{\frac{D_{{\rm op}}}{T}\log\left(\frac{T^{2}\Delta}{D_{{\rm op}}^{3}sr^{3/2}}\right),D_{\rm op}\right\}$. We have